OpenStudy (anonymous):

pleaseHELP SOLVE THIS: state the slope and y intercept of y=5/6x-3

OpenStudy (anonymous):

equations of the form y = mx + b have "m" as the slope and "b" as the y-intercept. You have an equation in that exact form with "m" = 5/6 and "b" = -3.

OpenStudy (anonymous):

So the answer is 5/6 is the slope and -3 is the intercept? I guess I have been overthinking it - don't I have to solve for X and Y?

OpenStudy (anonymous):

You got it! Good job! Yes, don't over-think.

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

You're welcome! If you like my work, click "best response"